[figure: 9.7 cm]

EVIDENCE FOR A MAGNETO-GALVANIC MECHANISM IN

ANISOTROPY OF RESISTIVITY IN U_{3}P_{4} AND U_{3}As_{4}

Piotr Wiśniewski and Zygmunt Henkie

Institute for Low Temperature and Structure Research, Polish Academy of Sciences, PO Box 1410, 50-950 Wrocław 2, Poland

U_{3}X_{4}-type compounds (X = P, As, Sb, Bi) are iso-structural (b.c.c.), multi-sublattice ferromagnets showing spectacular variety of magnetic and electronic properties due to the interplay of f-electron hybridization, crystal-field symmetry and magnetic interactions (m_{eff} up to ≈ 70m_{0}, Δρ/ρ up to 30%). Here we show how the angular dependence of thermoelectric power measured in low magnetic field can prove the role of magnetogalvanic effect in the anisotropy of their electrical resistivity. In both compounds resistivity displays very strong ferromagnetic anisotropy. Its value depends on angle between electric current and direction of magnetic moment in the sample. Moreover, for the magnetic field directed along [100]- or [110]-axes, sharp peaks of electric resistivity are observed. Our measurements of angular dependence of thermoelectric power show that the ferromagnetic anisotropy persists, whereas, the peaks at [100] and [110] do not appear. It proves that these peaks of the resistivity are of magneto-galvanic origin, being due to the electric current. We explain how these peaks may appear because of high Hall coefficient and domain structure of the samples.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Piotr Wiśniewski

Address for correspondence : Institute for Low Temperature and Structure Research, Polish Academy of Sciences, PO Box 1410, 50-950 Wrocław 2, Poland

Email address : wisniew@int.pan.wroc.pl